### v4

@@ Abstract
Forman-Ricci curvature (FRC) is a potent and powerful tool for analysing empirical networks, as the distribution of the curvature values can identify structural information that is not readily detected by other geometrical methods. Crucially, FRC captures higher-order structural information of clique complexes of a graph or Vietoris-Rips complexes, which is not readily accessible to alternative methods. However, existing FRC platforms are prohibitively computationally expensive. Therefore, herein we develop an efficient set-theoretic formulation for computing such high-order FRC in complex networks. Significantly, our set theory representation reveals previous computational bottlenecks and also accelerates the computation of FRC. Finally, We provide a pseudo-code, a software implementation coined FastForman, as well as a benchmark comparison with alternative implementations. We envisage that FastForman will be used in Topological and Geometrical Data analysis for high-dimensional complex data sets.
@@ 1 Introduction
Network analysis is one of the success stories of complex systems research. The basic is simple, to represent the network as a graph whose mathematical features can then be studied. Beyond the determination of clusters or the identification of particularly important vertices or edges, also certain statistical features can reveal valuable structural information. Among those, the so-called discrete Ricci curvatures (a name derived from their conceptual origin in Riemannian geometry) are particularly useful. And among those curvatures, the Forman Ricci curvature (FRC) is particularly easy to compute. But this comes at the expense of ignoring valuable information that is contained in higher-order relations between the vertices. To uncover such information, it is necessary to go to simplicial complexes, in the case at hand the clique complex of the graph. The topological aspects of such higher-order information are evaluated in Topological Data Analysis (TDA). There, with a scale parameter $r$, one constructs a graph by connecting pairs of data points whose distance is $\leq r$, and one then computes the homology, and in particular, the Betti numbers, of the resulting clique complex, called the Vietoris-Rips complex, as depending on $r$. This by now is a well-established method. But in line with the preceding, we want to extract geometric information, and as suggested, this should be done by the augmented FRC of such a complex.Now, these clique or Vietoris-Rips complexes are not arbitrary simplicial complexes but enjoy the special property that every simplex is automatically filled as soon as all its edges, that is, its 1-skeleton are present.However, while the FRC of a graph is very easy to compute, this is no longer so for the augmented FRC of a simplicial complex, at least with standard methods. Therefore, in this paper, we provide a systematic mathematical approach within which the FRC of such a Vietoris-Rips type complex can be easily and quickly computed. As the name indicates, FRC was originally introduced by Forman [12]. it was inspired by Bochner’s formula for Ricci tensors in Riemannian manifolds [1]. [12] provides a formula for FRC of a simplicial complex. The graph version has already found a wide range of applications. For instance, it has been applied in cancer diagnostics [20], detecting anomalies on brain networks [3], detection of dynamic changes on data sets [25], detection of stock market fragility [21] and to infer pandemic states [6].But in those and other cases, usually only the FRC of a graph was computed. To reveal the full power of the scheme, and to systematically compare the results with those obtained by TDA, we need to compute the FRC of such simplicial complexes. But in general, the time and memory complexity of computing higher-order dimensional information from complex networks is prohibitively expensive, which often limits applications. For example, the time for computing all cliques in a graph has exponential-time complexity [7, 11, 18]. Also, the Betti number computation required for TDA is an NP-hard problem [9], due to the matrix construction of boundary operators. Indeed, Betti number computation increases with the number of cliques and, consequently, requires memory-efficient algorithms. Likewise, the lack of efficient algorithms for computing generalized geometric structures has limited the use of high-order Forman-Ricci curvature (FRC) for complex networks. Set-theoretic approaches [10, 14, 28, 17] may offer a solution to this problem, and this is the approach of the present paper. To implement and apply this, we make an important observation that when the complex is of Vietoris-Rips type, that is, all simplices are filled when their 1-skeleton is present, this leads to considerable computational simplifications. And, incidentally, this is also a reason why TDA is such a powerful scheme in applications. Therefore, our assumption seems natural, and we then want to build an algorithm upon it. To this end, we developed a set-theoretic representation theory for higher-order network structures based on node neighbourhoods to give an alternative (and optimal) definition for FRC in terms of the classic node neighbourhood of a graph. This leads us to a representation theory that reveals previous computational bottlenecks and enables us to develop an efficient algorithm that markedly accelerates the computation of FRC in higher-order networks.We have implemented our algorithm - FastForman (see in [5]) - and benchmarked it with other implementations, namely GenerlisedFormanRicci [27] and HodgeLaplacians [23]. We found that our implementation boosted time processing and drastically reduced memory consumption during the process of computing FRC for low and high-dimensional network structures. These results reinforce our beliefs in the power of set-theoretic approaches to improve the efficiency of algorithm implementations for high-order geometric approaches to big data sets.
@@ 2 Network and higher-order network fundamentals
Let $x,y\in V$. We say that $y$ is a neighbour of $x$ if the edge $(x,y)\in E$. The neighbourhood of the node $x$ is defined by a set of nodes that are connected to $x$ via an edge in $E$, and we denote by $\pi_{x}$. Formally,
$\pi_{x}=\{y\in V\,|\,(x,y)\in E\},$ (1)
A simple example is shown in 2. We will define the Graph neighbourhood as the sequence of node neighbourhoods of a graph $G$. Formally,
@@
2.1 Abstract simplicial complex
The concepts of this section are as defined from [30, 29, 8]. Let $C$ be a set together with a collection $\mathcal{S}.$ We say that $C$ is an abstract simplicial complex if the following conditions are satisfied:
  1. 1.
    For each $v\in C$, $\{v\}\in\mathcal{S};$
  2. 2.
    If $\gamma\subseteq\alpha$ and $\alpha\in\mathcal{S}$, then $\gamma\in\mathcal{S}$.
We say that $\alpha$ are $d$-cells (or $(d+1)$-cliques) if $|\alpha|=d+1.$ When $\gamma\subseteq\alpha$, we say that $\gamma$ is a face of $\alpha$, or alternatively, $\alpha$ is a coface of $\gamma$. In our approach, in the special cases where $|\alpha|$=$|\gamma|$+1, we will denote $\gamma<\alpha$, or alternatively, $\alpha>\gamma$. We denote the set of all $d$-cells by $C_{d}$, and we can clearly reconstruct $C$ by taking $C=\bigcup\limits_{d}C_{d}.$ Let $G=(V,E)$ be a simple undirected graph. The collection $\mathcal{C}$ of all subsets of $\{x_{0},x_{1},\ldots,x_{d}\}\subseteq V$ such that the vertices span a cell of $C$ is called a vertex scheme of $C$. Clearly, $\mathcal{C}$ is a simplicial complex and can be identified to $C$ via isomorphism [16]. This result allows us to identify a cell $\alpha\in C_{d}$ with the a subset $\{x_{0},\ldots,x_{d}\}\subseteq V.$ Also, any combinations of subsets with size $k$ span a $(k-1)$-cell in $C_{k-1}$, for $k\in\{1,\ldots,d\}.$ We refer the reader to Fig. 1 and accompanying 3 for a graphic exposition of the concepts of simplicial complex and d-cells.
@@ 2.1 Abstract simplicial complex
$\pi_{\alpha}=\bigcap_{x\in\alpha}\pi_{x}.$ (3)
@@ 3 High-order Forman-Ricci curvature
Let $G=(V,E)$ a graph and $d\in\mathbb{N}\bigcup\{0\}$. Let $\alpha\in C_{d}$. We define the boundary of $\alpha$ by $\partial(\alpha)=\{\gamma\in C_{d-1},\,\gamma<\alpha\}$, for $d>0$ and $\partial(\alpha)=\emptyset$ for $d=0$. Alternatively, we can define $\partial(\alpha)=\{\{x_{0},\ldots,\hat{x_{i}},\ldots,x_{d}\},x_{i}\in\alpha\}=\{\gamma\subset\alpha\,,\,|\gamma|=d\}$, where $\hat{x_{i}}$ denotes the removal of element $x_{i}$. Note that $\partial(\alpha)$ has exactly $\binom{d+1}{d}=d+1$ elements. In 5 the concept of cell boundary is elucidated. We say that $\alpha_{1},\alpha_{2}\in C_{d}$ are neighbours if at least one of the following cell neighborhood conditions is reached:
@@
We say that $\alpha_{1}$ and $\alpha_{2}$ are parallel neighbours if conditions $1$ are reached, but not $2$ and we denote the parallelism by $\alpha_{1}//\alpha_{2}$. In case of both conditions, $1$ and $2$ are satisfied, $\alpha_{1}$ and $\alpha_{2}$ are said transverse neighbours. 6 elucidates the concept of cell neighbourhood. Let $\alpha\in C_{d}.$ We denote the set of all neighbours of a cell $\alpha$ by $N_{\alpha}$. The sets of parallel and transverse neighbours of $\alpha$ will be denoted by $P_{\alpha}$ and $T_{\alpha}$, respectively. We will also denote by $H_{\alpha}$ as the set of all $(d+1)$-cells that have $\alpha$ in its boundary, i.e, all $\beta\in C_{d+1}$, such that $\beta>\alpha$. From conditions $1$ and $2$, it is clear that these sets are disjoint, and we can also write
$N_{\alpha}=P_{\alpha}\sqcup T_{\alpha},$ (4)
@@
The $d$-th Forman-Ricci curvature (FRC) is then given as follows [26, 22]:
$\mathrm{F}_{d}(\alpha)=$ |$\left|\{\beta\in C_{d+1},\,\,\beta>\alpha\}\right|$ (5)
@@
$\mathrm{F}_{d}(\alpha)=\left|H_{\alpha}\right|+(d+1)-\left|P_{\alpha}\right|.$ (6)
@@
$\mathrm{F}(e)=\left|\{t\in C_{2},\,\,t>e\}\right|\,+2\,-\left|\{e^{\prime}\in E,\,\,e^{\prime}//e\}\right|,$ (7)
@@
$\mathrm{F}_{d}(C)=\dfrac{1}{|C_{d}|}\sum_{\alpha\in C_{d}}\mathrm{F}_{d}(\alpha).$ (8)
@@ 4.1 Set-theoretic representation of Forman-Ricci curvature
The set of neighbours of $\alpha$ can be redefined in terms of the node neighbourhood as follows (see Theorem 7.9):
$N_{\alpha}=\bigsqcup_{\gamma\in\partial(\alpha)}\,\bigcup_{\begin{subarray}{c}x\in\pi_{\gamma}\neq\emptyset\\ x\notin\alpha\end{subarray}}\{\gamma\cup\{x\}\},$ (11)
@@
$T_{\alpha}=\bigsqcup_{\gamma\in\partial(\alpha)}\,\bigcup_{\begin{subarray}{c}x\in\pi_{\gamma}\neq\emptyset\\ x\in\pi_{\alpha}\neq\emptyset\\ x\notin\alpha\end{subarray}}\{\gamma\cup\{x\}\}.$ (15)
From (15) and (4), we derive the set of parallel cells from a complementary set of transverse cells,
$P_{\alpha}=\bigsqcup_{\gamma\in\partial(\alpha)}\,\bigcup_{\begin{subarray}{c}x\in\pi_{\gamma}\neq\emptyset\\ x\notin\pi_{\alpha}\neq\emptyset\\ x\notin\alpha\end{subarray}}\{\gamma\cup\{x\}\}.$ (16)
In practical terms, equations (15) and (16) suggest that the transverse and parallel cells can be obtained from (11) by deciding if each neighbouring cell of $\alpha$ is contained in a $(d+1)$-cell or not. The details are shown in Theorem 7.15. Also, from Theorem 7.19,
$\left|T_{\alpha}\right|=(d+1)\left|H_{\alpha}\right|,$ (17)
which together with (6) provides
$\mathrm{F}(\alpha)=\frac{1}{(d+1)}|T_{\alpha}|+(d+1)-|P_{\alpha}|,$ (18)
which after expansion provides
$\mathrm{F}(\alpha)=\frac{1}{(d+1)}\sum_{\gamma\in\partial(\alpha)}|\pi_{\alpha}\cap\pi_{\gamma}-\alpha|+(d+1)-\sum_{\gamma\in\partial(\gamma)}|\pi_{\gamma}-\pi_{\alpha}-\alpha|.$ (19)
Moreover, we can express the number of cells of dimension $d+1$ containing $\alpha$ (see proof in Theorem 7.17) as
$H_{\alpha}=\bigcup_{x\in\pi_{\alpha}\neq\emptyset}\{\alpha\cup\{x\}\}.$ (20)
@@
$\mathrm{F}_{d}(\alpha)=|\pi_{\alpha}|+(d+1)-\sum_{\gamma\in\partial(\gamma)}|\pi_{\gamma}-\pi_{\alpha}-\alpha|.$ (21)
Equation (21) provides a direct computation of FRC from the parallel cells. However, an alternative equation can be derived by computing FRC directly from the cell neighbourhood. This is obtained by combining (4), (6) and Eq. 17 we have
@@
$\mathrm{F}_{d}(\alpha)=(d+2)\left|\pi_{\alpha}\right|+(d+1)-\left|N_{\alpha}\right|.$ (22)
The expansion of the term $\left|N_{\alpha}\right|$ leads to
$\left|N_{\alpha}\right|=\sum_{\gamma\in\partial(\alpha)}|\pi_{\gamma}-\alpha|=\sum_{\gamma\in\partial(\alpha)}|\pi_{\gamma}|-(d+1),$ (23)
@@
which when applied to (22) leads us to
@@
$\mathrm{F}_{d}(\alpha)=(d+2)\left|\pi_{\alpha}\right|+2\cdot(d+1)-\sum_{\gamma\in\partial(\alpha)}\left|\pi_{\gamma}\right|,$ (24)
Equations (19),(21) and (24) not only establish new formulations for FRC but also crucially enable us to propose novel algorithms to efficiently compute FRC in terms of the local node neighbourhoods of the cells into consideration. We propose Algorithms 4, 2, and 3 that should be parameterized by the graph $G=(V,E),$ the list of cells ($C$), and the maximum cell dimension ($d_{\max}$), and it computes FRC up to dimension $d\leq d_{\max}$.
Algorithm 2 Compute Forman-Ricci Curvature: Method A
@@
  for $\alpha\in C$ do
     $d:=|\alpha|-1$
     Compute $\mathrm{F}_{d}(c)$ according to (19):
     Compute $\pi_{\alpha};$
     for  $\gamma\in\partial(\alpha)$ do
        set $M:=\{v\in V\,|\,v\in\pi_{\gamma},\,v\notin\alpha\}$, $t:=0$, $p:=0;$
        for  $v\in M$ do
           if  $v\in\pi_{\alpha}$ then
              $t:=t+1$
           else
              $p:=p+1$
           end if
        end for
     end for
     set $\mathrm{F}_{d}(\alpha)=\frac{1}{(d+1)}\cdot t+(d+1)-p$
     Update $\mathcal{F}_{d}:=\mathcal{F}_{d}\cup\{\mathrm{F}_{d}(\alpha)\}$

### v5

@@ Abstract
Differential geometric approaches are ubiquitous in several fields of mathematics, physics and engineering, and their discretizations enable the development of network-based mathematical and computational frameworks, which are essential for large-scale data science. The Forman-Ricci curvature (FRC) - a statistical measure based on Riemannian geometry and designed for networks - is known for its high capacity for extracting geometric information from complex networks. However, extracting information from dense networks is still challenging due to the combinatorial explosion of high-order network structures. Motivated by this challenge we sought a set-theoretic representation theory for high-order network cells and FRC, as well as their associated concepts and properties, which together provide an alternative and efficient formulation for computing high-order FRC in complex networks. We provide a pseudo-code, a software implementation coined FastForman, as well as a benchmark comparison with alternative implementations. Crucially, our representation theory reveals previous computational bottlenecks and also accelerates the computation of FRC. As a consequence, our findings open new research possibilities in complex systems where higher-order geometric computations are required.
@@ 1 Introduction
Network analysis has emerged as a hot topic over the last decade [barabasi2016network, lewis2009network, borner2007network, Krioukov2012NetworkCosmology], and has enabled the theoretical and computational advances across several fields, for instance, in neuroscience [Stam2014ModernDisorders, Fan2017NetworkNino, Schoonheim2015NetworkSclerosis, Bassett2017NetworkNeuroscience, Derks2014NeuralPatients, gunnarcarlsson2018towards, santos2018topological, guidolin2022geometry, santos2023emergence, fulop2020we]. Topological Data Analysis (TDA) and Geometrical Data Analysis (GDA) emerged as alternative approaches for analyzing network-based data [Taylor2015TopologicalNetworks, zomorodian2012topological, pascucci2010topological, bubenik2015statistical]. Among many ways of extracting geometric information from complex networks, discrete Ricci curvatures are one of the most powerful approaches[samal2018comparative, forman2003bochner, Lin2011RicciGraphs]. Originally inspired by Bochner’s formula for Ricci tensors in compact manifolds [bochner1946vector], the discrete version of the Forman-Ricci curvature (FRC) was first introduced by the work of [forman2003bochner], where the authors provide a formula for FRC of a simplicial complex from the discrete Bochner and Laplacian operators. Its application is a wide range. For instance, it has been applied in cancer diagnostics [sandhu2015graph], detecting anomalies on brain networks [chatterjee2021detecting], detection of dynamic changes on data sets [weber2016forman], detection of stock market fragility [sandhu2016ricci] and to infer pandemic states [Barros_de_Souza_2021]. Despite the informational power of these techniques, the time and memory complexity of computing higher-order dimensional information from complex networks is prohibitively expensive, which prevents researchers from using existing algorithms for higher-order complex networks. For example, the time for computing all cliques in a graph has exponential-time complexity [dieu1986average, fellows2009clique, peeters2003maximum] and, when feasible, might require an efficient algorithm for this task. As another example, the Betti number computation [bochner1949curvature], due to the matrix construction of boundary operators is an NP-hard problem [edelsbrunner2014computational]. Indeed, Betti number computation increases with the number of cliques and, consequently, requires memory-efficient algorithms. Among the geometric approaches, the discrete Ricci curvature for low-dimensional structures is the most feasible. However, its computation only considers the interaction between nodes and edges, ignoring relevant structural information that might be contained in higher-order structures. In addition, the lack of efficient algorithms for computing generalized geometric structures limits the use of high-order Forman-Ricci curvature (FRC) for complex networks. Conveniently, one way to develop efficient and optimal algorithms is via set-theoretic approaches [enderton1977elements, jech2003set, zimmermann1985applications, pawlak2002rough]. To this end, we developed a set-theoretic representation theory for higher-order network structures based on node neighbourhoods to give an alternative (and optimal) definition for FRC in terms of the classic node neighbourhood of a graph. This leads us to a representation theory that reveals previous computational bottlenecks and enables us to develop an efficient algorithm that markedly accelerates the computation of FRC in higher-order networks. Lastly, we implemented our algorithm - FastForman (see in [kaggle_FRC_report]) - and benchmarked with other implementations, namely GenerlisedFormanRicci [generalisedformanricci] and HodgeLaplacians [hodgelaplacians]. Impressively, our implementation boosted time processing and drastically reduced memory consumption during the process of computing FRC for low and high-dimensional network structures. These results reinforce our beliefs in the power of set-theoretic approaches to improve the efficiency of algorithm implementations and their importance to motivate the work of scientists on applying high-order geometric approaches to big data sets.
@@ 2 Network and higher-order network fundamentals
Let $x,y\in V$. We say that $y$ is a neighbour of $x$ if the edge $(x,y)\in E$. The neighbourhood of the node $x$ is defined by
(1) $\pi_{x}=\{y\in V\,|\,x\sim y\},$
where $x\sim y$ denotes that $x$ is connected to $y$ via an edge in $E$. A simple example is shown in 2. We will define the Graph neighbourhood as the sequence of node neighbourhoods of a graph $G$. Formally,
@@
A subgraph $G^{{}^{\prime}}=(V^{{}^{\prime}},E^{{}^{\prime}})$ of $G$ (denoted by $G^{{}^{\prime}}\subseteq G$) is a graph such that $V^{{}^{\prime}}\subseteq V$ and $E^{\prime}=\{(x,y)\,|\,x,y\in V^{\prime},x\neq y\}$, which clearly satisfies $E^{{}^{\prime}}\subseteq E$. A complete subgraph in $G$ is a subgraph in which all nodes are pairwise connected by an edge. They are also denoted $cliques$ or $cells$ of a graph, which are concepts associated with higher-order networks since they represent indivisible objects in their own right.
Let $d\in\mathbb{N}\bigcup\{0\}$  then the complete subgraphs in $G$ with $d+1$ nodes are called (d+1)-vertex cliques or d-cells in the graph. In the present sequel, we only deal with complete subgraphs and, for convenience, we will refer to them simply as subgraphs. We will denote $C_{d}$ as the set of all $d$-cells in G. Given a $d$-cell, $\alpha\in C_{d},$ we can also explicitly define $\alpha=[x_{0},\ldots,x_{d}]$ such that $x_{i}\in V$ (with $i=0,\ldots,d$) and $\alpha$ is a complete subgraph of $G$, i.e., there exists $e\in E$ such that $e=(x_{i},x_{j}),$ for all $x_{i},x_{j}\in\alpha$, $x_{i}\neq x_{j}$. Under this formulation, a $d$-cell $\alpha=[x_{0},\dots,x_{d}]$ can also be seen as the complete subgraph $G_{\alpha}=(V_{\alpha},E_{\alpha})$ such that $V_{\alpha}=\{x_{0},\ldots,x_{d}\}$ and $E_{\alpha}=\{(x_{i},x_{j})\,|\,x_{i},x_{j}\in V_{\alpha},i\neq j\}.$ Herein we define a simplicial complex [edelsbrunner2022computational] as the union of all sets of complete subgraphs in $G$, i.e., $C=\bigcup\limits_{d}C_{d}.$ We refer the reader to Fig. 1 and accompanying 3 for a graphic exposition of the concepts of simplicial complex and d-cells.
@@
(3) $\pi_{\alpha}=\bigcap_{x\in\alpha}\pi_{x}.$
@@
Remark 2.1.
Note that the definition in (3) implies that if $x\in\pi_{\alpha}$ then $x\notin\alpha,$ since we are dealing with simple graphs.
Finally, we can characterize the set of nodes of $\alpha$ in terms of its node neighbourhood as
(4) $V_{\alpha}=\{x_{0},\ldots,x_{d}\}=\Bigg{\{}x\in\alpha\,|\,x\in\bigcap_{\begin{subarray}{c}y\in\alpha\\ y\neq x\end{subarray}}\pi_{y}\Bigg{\}}.$
This formulation enables us to identify $\alpha\sim V_{\alpha}$, which is $[x_{0},\ldots,x_{d}]\sim\{x_{0},\ldots,x_{d}\}$. As a consequence, it enables us to deal with complete subgraphs as their representation in terms of nodes and all the set operations can be applied in a natural manner. This will be crucial for defining the concepts of parallelism and transversality of cells in the following section.
@@ 3 High-order Forman-Ricci curvature
Let $G=(V,E)$ a graph and $d\in\mathbb{N}\bigcup\{0\}$. Let $\gamma\in C_{d-1}$ and $\alpha\in C_{d}$. We denote by $\partial(\alpha)$ the set of all complete subgraphs with $d$ nodes contained in $\alpha$, i.e., all $(d-1)$-cells contained in $\alpha$. Formally, $\partial(\alpha)=\{\gamma\in C_{d-1},\,\gamma\subset\alpha\}$, for $d>0$ and $\partial(\alpha)=\emptyset$ for $d=0$. In analogy with our set-theoretic approach we alternatively define $\partial(\alpha)=\{\{x_{0},\ldots,\hat{x_{i}},\ldots,x_{d}\},x_{i}\in\alpha\}=\{\gamma\subset\alpha\,,\,|\gamma|=d\}$, where $\hat{x_{i}}$ denotes the removal of element $x_{i}$. Note that $\partial(\alpha)$ has exactly $\binom{d+1}{d}=d+1$ elements. In 5 the concept of cell boundary is elucidated. We denote that $\gamma$ is contained in the boundary of $\alpha$ by $\gamma<\alpha$ (noteworthy, in terms of set properties, it also implies that $\gamma\subset\alpha$). Analogously, we denote that $\alpha$ contains $\gamma$ in its boundary by $\alpha>\gamma$ (that is, in set theory formulation, $\alpha\supset\gamma$). We say that $\alpha_{1},\alpha_{2}\in C_{d}$ are neighbours if at least one of the following cell neighborhood conditions is reached:
@@
We say that $\alpha_{1}$ and $\alpha_{2}$ are parallel neighbours if conditions $1$ is reached, but not $2$ and we denote the parallelism by $\alpha_{1}//\alpha_{2}$. In case of both conditions, $1$ and $2$ are satisfied, $\alpha_{1}$ and $\alpha_{2}$ are said transverse neighbours. 6 elucidates the concept of cell neighbourhood. Let $\alpha\in C_{d}.$ We denote the set of all neighbours of a cell $\alpha$ by $N^{d}_{\alpha}$. The sets of parallel and transverse neighbours of $\alpha$ will be denoted by $P^{d}_{\alpha}$ and $T^{d}_{\alpha}$, respectively. We will also denote by $H^{d}_{\alpha}$ as the set of all $(d+1)$-cells that have $\alpha$ in its boundary, i.e, all $\beta\in C_{d+1}$, such that $\beta>\alpha$. From conditions $1$ and $2$, it is clear that these sets are disjoint, and we can also write
(5) $N^{d}_{\alpha}=P^{d}_{\alpha}\sqcup T^{d}_{\alpha},$
@@
The $d$-th Forman-Ricci curvature (FRC) is then given as follows [weber2017characterizing, saucan2018forman]:
(6) $\mathrm{F}_{d}(\alpha)=$ |$\left|\{\beta\in C_{d+1},\,\,\beta>\alpha\}\right|$
@@
(7) $\mathrm{F}_{d}(\alpha)=\left|H^{d}_{\alpha}\right|+(d+1)-\left|P^{d}_{\alpha}\right|.$
@@
(8) $\mathrm{F}(e)=\left|\{t\in C_{2},\,\,t>e\}\right|\,+2\,-\left|\{e^{\prime}\in E,\,\,e^{\prime}//e\}\right|,$
@@
(9) $\mathrm{F}_{d}(C)=\dfrac{1}{|C_{d}|}\sum_{\alpha\in C_{d}}\mathrm{F}_{d}(\alpha).$
@@ 4.1 Set-theoretic representation of Forman-Ricci curvature
First, note that we can formulate the set of $d$-cells as
(10) $C_{d}=\Big{\{}\alpha:=\{x_{0},\ldots,x_{d}\},\,\,x_{i}\in\pi_{\alpha\setminus\{x_{i}\}},\,\forall i\Big{\}}.$
The details are shown in Theorem B.1. The set of neighbours of $\alpha$ can be redefined in terms of the node neighbourhood as follows (see Theorem B.9):
(11) $N^{d}_{\alpha}=\bigsqcup_{\gamma\in\partial(\alpha)}\,\bigcup_{x\in\pi_{\gamma}\setminus(\alpha\setminus\gamma)\neq\emptyset}\{\gamma\cup\{x\}\},$
@@
(12) $T^{d}_{\alpha}=\bigsqcup_{\gamma\in\partial(\alpha)}\,\bigcup_{x\in\pi_{\gamma}\cap\pi_{\alpha\setminus\gamma}\neq\emptyset}\{\gamma\cup\{x\}\}.$
Remark 4.1.
Indeed, we have that $\pi_{\alpha}=\pi_{\gamma}\cap\pi_{\alpha\setminus\gamma,\,\forall\gamma\in\partial(\alpha)}$. However, we emphasize the right-hand side of this expansion in equations (11) and (12) in order to visualize (or intuitively conceptualise) the construction of those sets by cell boundary.
Moreover, we can express the number of cells of dimension $d+1$ containing $\alpha$ (see proof in Theorem B.13) as
(13) $H^{d}_{\alpha}=\bigcup_{x\in\pi_{\alpha}\neq\emptyset}\{\alpha\cup\{x\}\}.$
@@
(14) $P^{d}_{\alpha}=\bigsqcup_{\gamma\in\partial(\alpha)}\bigcup_{\begin{subarray}{c}\Pi^{\gamma}_{\alpha}\neq\emptyset\\ x\in\Pi^{\gamma}_{\alpha}\end{subarray}}\{\gamma\cup\{x\}\},$
where
(15) $\Pi^{\gamma}_{\alpha}=\pi_{\gamma}\setminus(\pi_{\alpha}\cup(\alpha\setminus\gamma)).$
The details are shown in Theorem B.17. Combining equations (7), (13) and (14), and considering $\alpha=\{x_{0},\ldots,x_{d}\}$ we finally have that
@@
(16) $\mathrm{F}_{d}(\alpha)=\left|\bigcap\limits^{d}_{i=0}\pi_{x_{i}}\right|\,+(d+1)\,-\left|\bigsqcup_{i=0}^{d}\bigcap\limits^{d}_{\begin{subarray}{c}j=0\\ j\neq i\end{subarray}}\pi_{x_{j}}\setminus{(\pi_{x_{i}}}\cup\{x_{i}\})\right|.$
Equation (16) provides a direct computation of FRC from the parallel cells. However, an alternative equation can be derived by computing FRC directly from the cell neighbourhood This is obtained by combining (5) and (7), and using the fact that (Theorem B.15)
(17) $\left|T^{d}_{\alpha}\right|=(d+1)\left|H^{d}_{\alpha}\right|,$
which results explicitly in
(18) $\mathrm{F}_{d}(\alpha)=(d+2)\left|\pi_{\alpha}\right|+(d+1)-\left|N^{d}_{\alpha}\right|.$
The expansion of the term $\left|N^{d}_{\alpha}\right|$ leads to
(19) $\left|N^{d}_{\alpha}\right|=\sum_{i=0}^{d}\left|\bigcap_{\begin{subarray}{c}j=0\\ j\neq i\end{subarray}}^{d}\pi_{x_{j}}-\{x_{i}\}\right|=\sum_{i=0}^{d}\left|\bigcap_{\begin{subarray}{c}j=0\\ j\neq i\end{subarray}}^{d}\pi_{x_{j}}\right|-(d+1),$
@@
which when applied to (18) and further simplification, finally leads us to our principle equation
@@
(20) $\mathrm{F}_{d}(\alpha)=(d+2)\left|\bigcap_{i=0}^{d}{}\pi_{x_{i}}\right|+2\cdot(d+1)-\sum_{i=0}^{d}\left|\bigcap_{\begin{subarray}{c}j=0\\ j\neq i\end{subarray}}^{d}\pi_{x_{j}}\right|.$
The details for reaching the formula in (20) is shown in Theorem B.19. Equation (20) not only establishes a new formulation for FRC but also crucially enables us to propose a novel algorithm to efficiently compute FRC in terms of the local node neighbourhoods of the cells into consideration. Indeed, we propose Algorithm 2 that should be parameterized by the graph $G=(V,E),$ the list of cells ($C$), and the maximum cell dimension ($d_{\max}$), and it computes FRC up to dimension $d\leq d_{\max}$. We subsequently address the practical implementation of equation (20). Assume $\alpha\in C_{d}$; then (20) computes the higher-order cells containing $\alpha$ by counting the nodes that are in the intersection of neighbouring nodes of $\alpha$ (i.e., $|\pi_{\alpha}|$), which are the nodes of the $(d+1)$-cells that contain $\alpha$ in their boundary. Moreover, the neighbouring cells of $\alpha$ are counted by the $d$-cells that have common boundaries with $\alpha$. By counting the number of neighbouring nodes that have these common boundaries (except the nodes in $\alpha$ itself, in order to avoid duplicates) we count all the neighbours of $\alpha$. In the simplest case, for $d=1$, to compute the FRC of an edge $\alpha=(x,y)$ (i.e. $F_{1}(\alpha)$), we determine the number of triangles (i.e. $2$-cells) containing the edge $\alpha$. This is achieved by computing the number of elements resulting from the intersection of the nodes that are neighbours of $x$ and $y$. Specifically, the number of neighbours of $\alpha$ is the result of counting the number of elements in the neighbourhood of $x$ and respectively, the neighbourhood of $y$, except the nodes $y$ and $x$ themselves. Analogously, for $d=2$, to compute FRC of a triangle $\alpha=\{x,y,z\}$ (i.e. $F_{2}(\alpha)$), the number of elements at the intersection between the neighbourhood of $x$, $y$ and $z$ provides the number of tetrahedrons (i.e. $3$-cells) containing $\alpha$. In other words, suppose we have the following intersection: $\pi_{x}\cap\pi_{y}\cap\pi_{z}=\{w_{1},...,w_{k}\}$. Then this intersection provides the set of nodes that together with the nodes of the triangle $\{x,y,z\}$ are the tetrahedrons that contain $\alpha$, that is, $\{\{w_{1},x,y,z\},\{w_{2},x,y,z\},...,\{w_{k},x,y,z\}\}=\{\alpha\cup\{w_{1}\},\ldots,\alpha\cup\{w_{k}\}\}$. Noteworthy, the number of neighbours of $\alpha$ are computed by counting the nodes that have $\{x,y\}$, $\{x,z\}$ and $\{y,z\}$ as common boundary and are not in $\alpha$. To facilitate intuition about these computations, we provide a graphical representation of the resulting computations for the case of $2$-cells in Fig. 2.
Figure 2: Example of how the algorithm detects cell neighbours from node neighbourhood for $2$-cells (triangles). The $2$-cell $\alpha$ is enhanced in orange, whilst its boundary is drawn in red. According to the algorithm, the number of $3$-cells containing $\alpha$ in its boundary ($|H^{d}_{\alpha}|$) coincides with the number of common neighbours of the nodes in $\alpha$ ($|\pi_{\alpha}|$). Also, the number of neighbours per boundary is counted by the size of intersections of node neighbourhoods in the boundary. In this example, $\mathrm{F}_{2}(\alpha)=4\cdot|\{6\}|+3-(|\{3,6\}|+|\{6\}|+|\{6\}|)=3.$ The same output is reached by using (6): $\mathrm{F}_{2}(\alpha)=1+3-1=3.$
Algorithm 2 Compute Forman-Ricci Curvature: FastForman
@@
  for $c\in C$ do
     $d:=|c|-1$
     Compute $\mathrm{F}_{d}(c)$ according to (20)
     Update $\mathcal{F}_{d}:=\mathcal{F}_{d}\cup\{\mathrm{F}_{d}(c)\}$